OpenStudy (anonymous):

need help Determine the area required for a new landfill site with a projected life of 30 years for a population of 250,000 generating 2.02 kg per person per day. The density of the compacted waste is 470 kg /m3. The height of the landfill is not to exceed 15 m. Assume the land fill will be a rectangular cube. @RoseDryer this is the only problem left i need you to check for me. I'll tell you what i get for each and let me know if i'm correct please.

OpenStudy (mathstudent55):

Start with the population and multiply by the amount of waste produced per person per day. That gives you the total amount of waste per day in kg.

OpenStudy (mathstudent55):

Then multiply that amount by 365.25 to have the amount of waste in 1 year. Then multiply the last amount by 30 to find the total amount of waste produced in 30 years in kg.

OpenStudy (mathstudent55):

Are you following this?

OpenStudy (anonymous):

over 30 year population i get 250000*30=7500000 per day i get 250000*2.02=505000 per year i get 250000*2.02*365 184325000 total wast in 30 years i get 250000*2.02*365*30= 5529750000 then i do 5529750000/470=11765425.53 and theni divided 11765425.53/15=784361.702 are these number correct?

OpenStudy (mathstudent55):

Yes. Everything is correct. Using your numbers, the answer is 784361 m^2. The only thing I would have done differently is I'd use 365.25 days per year. Since every 4 years, there is a leap year with 366 days, the avrage number of days in a year is 365.25, not 365.

OpenStudy (mathstudent55):

\(\dfrac{2.02 ~kg}{person - day} \times 250,000 ~persons \times \dfrac{365.25 ~days}{year} \times 30 ~years \times \dfrac{m^3}{470~kg} \) \(=\dfrac{2.02 ~\cancel{kg}}{\cancel{person} - \cancel{day}} \times 250,000 ~\cancel{persons} \times \dfrac{365.25 ~\cancel{days}}{\cancel{year}} \times 30 ~\cancel{years} \times \dfrac{m^3}{470~\cancel{kg}} \) \(= 11~773~484 ~m^3\)

OpenStudy (anonymous):

oh okay, thanks